# Same Task, Different Circuits: Disentangling Modality-Specific Mechanisms in VLMs

### v2

@@ 5 Improving Visual Tasks Performance
We find that alignment between images and analogous texts gradually increases across layers (Figure 7). However, increase is slow, with higher similarities only occurring later in the model, after data token activations have been moved to later positions (this information flow was shown in Kaduri et al. (2024), and is seen in Figure 3). We hypothesize that if these text-aligned representations were available earlier in the model’s processing, it could improve performance on visual tasks.
@@
Table 1: Back-patching visual token embeddings increases accuracy across tasks. We measure the accuracy improvement of VLMs on our visual tasks, while back-patching visual token embeddings from later layers back to earlier layers. This method increases accuracy at varying levels, bridging an average of 32% of the gap between VLM accuracy for visual and textual tasks.
Model Counting Arithmetic Color Ordering Factual Recall Sentiment Analysis
Qwen2-7B-VL $+4.2\%$ $+9.4\%$ $+0.6\%$ $+0.3\%$ $+2.1\%$
Pixtral-12B $+1.5\%$ $+7.6\%$ $+4.0\%$ $+2.4\%$ $+17.8\%$
Gemma-3-12B $+2.9\%$ $+6.7\%$ $+1.5\%$ $+1.8\%$ $+5.5\%$
To utilize the more textually-aligned representations of visual tokens, we employ back-patching (Biran et al., 2024; Lepori et al., 2024): the process of patching the late representations back to earlier layers (Figure 1c). For each visual prompt, we first perform a forward pass through the model and record activations $a^{l}$ for a window of consecutive layers centered at $l_{src}$, specifically $l\in\{l_{src}-i,\ldots,l_{src},\ldots,l_{src}+i\}$ where $i\in\{0,1,2\}$ determines the window size (1, 3, or 5 layers). We then perform another forward pass, but this time replace the activations at a corresponding window centered at $l_{dst}$ (where $l_{dst}<l_{src}$) with our previously recorded activations from around $l_{src}$. This replacement is performed in parallel across all visual token positions. In a model with $L$ layers, we systematically explore this back-patching by varying the source layers $l_{src}\in\{\lfloor\frac{L}{2}\rfloor+1,\ldots,L-1\}$ and destination layers $l_{dst}\in\{0,1,\ldots,\lfloor\frac{L}{2}\rfloor\}$, while maintaining the same window size for both source and destination, and choose the best performing $l_{src},l_{dst}$ (reported in Section E.3).
When applying back-patching, the results (Table 1) show an average absolute improvement in accuracy of $4.6\%$ on visual prompts, closing approximately $32\%$ of the performance gap between visual and textual accuracy. Additional back-patching experiments are detailed in Section E.1 and Section E.2. To confirm that these improvements stem from leveraging textually-aligned representations rather than mere additional computation, we conduct a control experiment and apply similar back-patching to textual prompt variants. This control yields smaller accuracy gains in roughly 79% of cases (further shown in Section E.4), supporting our hypothesis that improved visual performance is more due to the early transition of visual tokens to textually-aligned representations.
@@ 7 Conclusions, Limitations and Discussion
We focused on question-answering tasks involving single images, across varied domains. Further work could explore multi-image scenarios or more complex and visually-inclined reasoning tasks, where the model is expected to perform better on visual representations, such as navigating through visual environments. We limited our scope to simpler tasks as current open models struggle with more complex visual reasoning, making mechanism analysis less meaningful in these cases, which are left for future work as models advance.
Our experiments show that patching visual token representations from later to early layers closes 32% of the performance gap between analogous visual and textual prompts. Yet, the causes for the remaining gap remain elusive, suggesting potential for further visual processing improvements. For instance, the limited benefit of back-patching in visual factual recall indicates that enhanced processing in the query positions may be needed to better identify visual entities.
@@ Acknowledgments
We thank Tal Haklay for her feedback on this project, as well as Simon Schrodi, Jiahai Feng, and Tomer Ashuach for their helpful comments on drafts of this paper. This research was supported by an Azrieli Foundation Early Career Faculty Fellowship and Open Philanthropy. DA is supported by the Ariane de Rothschild Women Doctoral Program. YG is supported by the Google Fellowship. This research was funded by the European Union (ERC, Control-LM, 101165402). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Definition of MLP Neurons
@@
$\mathbf{h}_{post}^{l}=\sigma(\mathbf{h}_{in}^{l}\mathbf{W}_{gate}^{{l}^{\top}})\circ(\mathbf{h}_{in}^{l}\mathbf{W}_{in}^{{l}^{\top}})$ (6)
$\mathbf{h}_{out}^{l}=\mathbf{h}_{post}^{l}\mathbf{W}_{out}^{l}$ (7)
@@ B.1 Prompt examples
Table 2 presents prompt examples for each task in both input modalities.
@@
The amount of prompts per task is shown in Table 3. The amount of prompts for the same task and modality varies between models due to tokenization considerations (e.g. all object colors in the spatial ordering task must be tokenized to a single token for positional alignment purposes—prompts that contain multi-token object colors are filtered out).
@@ B.2 Tasks Generation Details
For the textual variant, each prompt consists of two two-digit operands, ensuring positional alignment between different prompts (e.g., the second position always includes the singles digit of the first operand). This consistency is important because all analyzed models tokenize numbers into separate digits. For the visual variant, we create white-background images ($75$×$338$ resolution) displaying the identical arithmetic calculations in large black font centered in the image, maintaining a simple presentation that focuses entirely on the calculation itself.
@@
For the textual variant, we select entities (to replace the [X] in the prompt, e.g. “Michael Jordan”) that, when combined with the prefix “Consider [X]”, result in exactly 5 tokens to ensure positional alignment. We use this number as it results in the highest number of possible entities in the given dataset. For the visual variant, we collect entity images with public licenses from Wikimedia Commons. We rank 10 potential images per entity using CLIP (Radford et al., 2021) scores and select the best match for each entity. We resize and center-crop each image (to $256$x256256256256 resolution). We manually review the images to remove any that contain “shortcut” hints (e.g., any image of a basketball player holding a basketball), to ensure models rely on factual recall rather than visual cues. In both variants, the counterfactual prompts for each template are sampled only from within the same template.
@@ B.3 Circuit Discovery and Evaluation
For our circuit discovery experiments (detailed in Section 2 and Section 4), we allocate each modality-specific task dataset with a 75/25 split: 75% of the prompts for discovery and 25% of the prompts for faithfulness evaluation. We divide prompts between the discovery and evaluation subsets such that all possible answers are distributed in the same ratio between the subsets.
When answers span multiple tokens (e.g. in the arithmetic task, where answers can be comprised of multiple digits), we measure the patching effect (Equation 1) solely on the first token. We verify that the answers for each prompt and counterfactual prompt ($r$ and $r^{\prime}$) differ in the analyzed first token.
@@ B.4 Compute Resources
Our experiments were conducted on an NVIDIA L40404040 node equipped with $8$ GPUs, each containing $48$GB of memory. Peak memory consumption occurred during circuit discovery operations on the Gemma-3-12B-Instruct model, that required the parallel use of $4$ GPUs. The complete experimental suite, including studies not featured in the final paper, consumed roughly 200–300 GPU hours.
@@ Appendix C Task Accuracies
In Table 4 we report the accuracies of models on each of the tasks. We observe that across most tasks, VLMs achieve higher accuracy in the textual variant than the visual variant, motivating our analysis into this performance gap. This is compatible with earlier results on closed-source VLMs (Fu et al., 2024).
@@ Appendix D Additional Circuit Discovery Results
In Table 5 we report the circuit size, in a percentage of the model’s component count. As described in Section 4, the circuits are chosen to be the minimum-sized circuit with a faithfulness of over 80%. The faithfulness achieved by each circuit on each task and modality is reported in Table 6.
@@
Furthermore, in Figure 10, Figure 11, Figure 12, Figure 13, Figure 14 we present the patching importance scores for each model and task, summed per position and layer. These figures are complementary to Figure 3.
@@ E.1 Iterative Back-patching
Extending the back-patching experiments in Section 5, we explore if additional processing of visual data tokens yields further improvements. To do that, we perform back-patching iteratively (back-patching more than once). After identifying the optimal layers $l_{src},l_{dst}$ and layer window size for each model and task, we perform back-patching iteratively. Namely, after each patching of the window centered at $l_{dst}$, we continue the forward pass to calculate the window centered at $l_{src}$, and re-patch the window centered at $l_{dst}$. This process is repeated up to 10 times. As shown in Figure 8, performance degrades after the first back-patching iteration across all but one task and model pairs, with each subsequent iteration further reducing model accuracy. We attribute this decline to representations becoming increasingly out-of-distribution with each iteration, making them progressively less compatible with the model’s learned parameters.
@@
E.2 Back-Patching in VQAv2
To demonstrate that back-patching generalizes beyond our constructed dataset in broader visual question-answering, we evaluate its impact on model accuracy using the standard VQAv2 dataset (Goyal et al., 2017). We measure the accuracy of each analyzed VLM on $3000$ randomly sampled visual prompts (ensuring no image repetition) as baseline performance, and measure it again when applying back-patching (as described in Section 5). We limit the analysis to $3000$ random prompts due to computational requirements of exploring different back-patching settings. The results confirm the average improvements observed in Section 5: back-patching enhances visual accuracy by 3.2% for Qwen2-7B-VL ($65.0\%\to 68.2\%$), 4.0% for Pixtral-12B ($62.2\%\to 66.2\%$), and 4.6% for Gemma-3-12B ($65.7\%\to 70.3\%$). Since the visual prompts in VQAv2 do not have textual analogs, we do not perform the control experiment in this setting.
E.3 Best Layers For Back-Patching
For each model and task, Table 7 shows the values of $l_{src},l_{dst}$ and the layer window size that lead to the highest back-patching results, reported in Table 1.
@@ E.3 Best Layers For Back-Patching
These optimal layers match our hypothesis across models: replacing visual data token representations at model layers when they are less textually-aligned, with more textually-aligned representations, yields the highest accuracy improvements. This pattern persists even in Gemma-3-12B, where visual data tokens demonstrate high alignment with textual analogs throughout most of the model (as shown in Figure 7): tasks showing the greatest accuracy gains from back-patching (counting, arithmetic, and sentiment analysis, as detailed in Table 1) benefit from patching in the earliest layers where the alignment of tokens with textual analogs is relatively low.
E.4 Back-Patching Control Experiments
@@
This section presents our back-patching control experiment results across models and tasks. The accuracy increases from back-patching visual data tokens (reported in Table 1) can either stem from earlier textually-aligned representations or from additional compute. To isolate the effect of adding computational layers, we conduct the same back-patching analysis on analogous textual prompts. In this control experiment, the alignment with textual tokens is inherent, so any accuracy increase must stem from added compute. Across each back-patching setting—model, task and layer window size—we check for how many values of $l_{src}$ and $l_{dst}$ is the accuracy increase caused by back-patching larger than the control accuracy increase. Our findings (Figure 9) show that in most settings, visual back-patching yields greater accuracy improvements than the control for a majority of layer values. A notable exception occurs in factual recall tasks for Qwen2-VL-7B and Gemma-3-12B, where back-patching shows minimal accuracy gains and is outperformed by the control in over half the cases. These results support our hypothesis that improved visual performance stems primarily from the early transition of visual tokens into textually-aligned representations, rather than from increased computational depth.

### v3

@@ 5 Improving Visual Tasks Performance
We find that alignment between images and analogous texts gradually increases across layers (Figure˜7). However, this increase is slow, with higher similarities only occurring later in the model, after information from data tokens moves to later positions (this information flow was shown in Kaduri et al. (2024), and is seen in Figure˜3). We hypothesize that if these text-aligned representations were available earlier in the model’s processing, it could improve performance on visual tasks.
@@
Table 1: Back-patching visual token embeddings increases accuracy across tasks. We measure the improvement on visual tasks while back-patching visual token embeddings from later layers back to earlier layers. We report the back-patched accuracies and standard deviations, as well as the change from baseline model performance. Green marks statistically significant improvement. Back-patching shows an increase from baseline accuracies (reported in Appendix˜C) across most settings.
Task Qwen2-7B-VL Pixtral-12B Gemma-3-12B
Counting $75.0\%$ ($4.2\%$ $\pm$ $1.9\%$) $68.3\%$ ($2.2\%$ $\pm$ $1.1\%$) $75.4\%$ ($2.0\%$ $\pm$ $2.1\%$)
Arithmetic $76.9\%$ ($9.4\%$ $\pm$ $1.3\%$) $29.4\%$ ($6.7\%$ $\pm$ $1.6\%$) $94.2\%$ ($6.7\%$ $\pm$ $0.7\%$)
Color Ordering $74.8\%$ ($\% ±\pm 0.60.6\%$) $84.7\%$ ($\% ±\pm 1.41.4\%$) $47.5\%$ ($1.5\%$ $\pm$ $0.9\%$)
Factual Recall $69.7\%$ ($1.6\%$ $\pm$ $2.2\%$) $49.1\%$ ($4.1\%$ $\pm$ $2.3\%$) $68.6\%$ ($2.1\%$ $\pm$ $1.8\%$)
Sentiment Analysis $94.8\%$ ($2.2\%$ $\pm$ $1.4\%$) $87.4\%$ ($17.7\%$ $\pm$ $2.2\%$) $98.0\%$ ($5.4\%$ $\pm$ $0.9\%$)
To utilize the more textually-aligned representations of visual tokens, we employ back-patching (Biran et al., 2024; Lepori et al., 2024): the process of patching the late representations back to earlier layers (Figure˜1c). For each visual prompt, we first perform a forward pass through the model and record activations $a^{l}$ for a window of consecutive layers centered at $l_{src}$, specifically $l\in\{l_{src}-i,\ldots,l_{src},\ldots,l_{src}+i\}$ where $i\in\{0,1,2\}$ determines the window size (1, 3, or 5 layers). We then perform another forward pass, but this time replace the activations at a corresponding window centered at $l_{dst}$ (where $l_{dst}<l_{src}$) with our previously recorded activations from around $l_{src}$. This replacement is performed in parallel across all visual token positions. In a model with $L$ layers, we systematically explore this back-patching by varying the source layers $l_{src}\in\{\lfloor\frac{L}{2}\rfloor+1,\ldots,L-1\}$ and destination layers $l_{dst}\in\{0,1,\ldots,\lfloor\frac{L}{2}\rfloor\}$, while maintaining the same window size for both source and destination, and choose the best performing $l_{src},l_{dst}$ (reported in Section˜E.2).
To show statistical significance, we use bootstrap re-sampling with $1000$ iterations and verify that the lower bound of the back-patched accuracy is higher than the baseline accuracy. We report the back-patched accuracies, the improvement compared to the baseline model accuracy, and standard deviations across resamples in Table˜1.
Our results show that back-patching leads to an average absolute improvement of $4.6\%$ on visual prompts, closing approximately $32\%$ of the performance gap between visual and textual accuracy. Additional back-patching experiments are detailed in Section˜E.1 and Section˜E.4.
To confirm that these improvements stem from leveraging textually-aligned representations rather than mere additional computation, we conduct a control experiment and apply similar back-patching to textual prompt variants. This control yields smaller accuracy gains in roughly 79% of cases (further shown in Section˜E.3), supporting our hypothesis that improved visual performance is more due to the early transition of visual tokens to textually-aligned representations.
@@ 7 Conclusions
8 Limitations
We focused on question-answering tasks involving single images, where the image appears prior to the query. Further work could explore multi-image scenarios, different prompt order, or more complex and visually-inclined reasoning tasks, where the model is expected to perform better on visual representations (e.g., navigating through a visual environment). We limited our scope to simpler tasks with single-token completions due to two key constraints. First, current open models struggle with more complex visual reasoning, making mechanism analysis less meaningful in these cases, and second, current circuit analysis methods can only be applied to single-token completions.
Our experiments show that patching visual token representations from later to early layers closes 32% of the performance gap between analogous visual and textual prompts. Yet, the causes for the remaining gap remain unclear, suggesting potential for further visual processing improvements.
Finally, we focus our analysis on adapter-based VLMs, as this is the current state-of-the-art approach for aligning vision and language. Our findings may not immediately apply to models utilizing different fusing approaches, which we leave for future work.
@@ Acknowledgments
We thank Tal Haklay for her feedback on this project, as well as Simon Schrodi, Jiahai Feng, and Tomer Ashuach for their helpful comments on drafts of this paper. We thank the anonymous reviewers of this paper for providing useful feedback. This research was supported by an Azrieli Foundation Early Career Faculty Fellowship and Open Philanthropy. DA is supported by the Ariane de Rothschild Women Doctoral Program. YG is supported by the Google Fellowship. This research was funded by the European Union (ERC, Control-LM, 101165402). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Formal circuit definitions
A.1 Definition of MLP Neurons
@@
$\mathbf{h}_{post}^{l}=\sigma(\mathbf{h}_{in}^{l}\mathbf{W}_{gate}^{{l}^{\top}})\circ(\mathbf{h}_{in}^{l}\mathbf{W}_{in}^{{l}^{\top}})\ ,$ (6)
$\mathbf{h}_{out}^{l}=\mathbf{h}_{post}^{l}\mathbf{W}_{out}^{l}\ ,$ (7)
@@
A.2 Definition of circuits
A circuit $\mathbf{c}$ is formally defined as a set of components:
$\mathbf{c}=\{(l,p,i)\mid l\in\mathbb{N},p\in\mathbb{N},i\in\mathbb{N}\}$
where each component is represented by a 3-tuple $(l,p,i)$ denoting the layer, position, and index respectively. Each component represents either an attention head (where $i$ denotes the attention head index within layer $l$) or an MLP neuron (where $i$ denotes the neuron index within layer $l$, as defined in Section˜A.1).
When decomposing a circuit into its data ($\mathbf{c}^{D}$), query ($\mathbf{c}^{Q}$), and generation ($\mathbf{c}^{G}$) sub-circuits, the circuit components are partitioned according to their position indices, such that $\mathbf{c}=\mathbf{c}^{D}\cup\mathbf{c}^{Q}\cup\mathbf{c}^{G}$. Formally, this decomposition is defined as:
$\mathbf{c}^{D}=\{(l,p,i)\in\mathbf{c}\mid p\in P_{D}\}$
$\mathbf{c}^{Q}=\{(l,p,i)\in\mathbf{c}\mid p\in P_{Q}\}$
$\mathbf{c}^{G}=\{(l,p,i)\in\mathbf{c}\mid p\in P_{G}\}$
where $P_{D}$, $P_{Q}$, and $P_{G}$ are disjoint sets of position indices corresponding to the data, query, and generation regions respectively. In our case, the generation region includes only the last input position.
A.3 Definition of circuit component mapping
A positional mapping is a function $M:\mathbb{N}\rightarrow\mathcal{P}(\mathbb{N})$ that maps each position in a textual task prompt to one or more positions in a visual task prompt containing equivalent information, and vice versa to map visual positions to textual positions (marked $M^{-1}$). We manually define the positional mapping for each task and model combination. For an individual circuit component $(l,p,i)$, the mapping transforms it to a set of components in the target modality:
$M\left(\left(l,p,i\right)\right)=\{(l,p^{\prime},i)\mid p^{\prime}\in M(p)\}\ .$
For example, in the object counting task with Qwen2-7B-VL, the mapping function $M$ maps textual position $23$ to visual position $96$, where both contain the first token of the query (“How”). The complete positional mappings for all tasks and models are provided in the code accompanying the paper.
To compute our similarity metrics (Equation˜4, Equation˜5) over query and generation sub-circuits, we map a sub-circuit from one modality to another (e.g. $\mathbf{c}^{Q}_{L\to V}$) by applying the mapping function to each component individually. Formally, for a query ($\mathbf{c}^{Q}$) or generation sub-circuit ($\mathbf{c}^{G}$), we obtain the mapped sub-circuit:
$\mathbf{c}_{L\to V}=M(\mathbf{c})=\bigcup_{(l,p,i)\in\mathbf{c}}M((l,p,i))\ .$
The treatment of data sub-circuits differs because a positional correspondence does not always exist between modalities in the data positions. For instance, in the sentiment analysis task, general words in the textual sequence may not map to specific visual tokens, but rather convey an overall sentiment or atmosphere that is distributed across the entire visual input. Therefore, when computing similarity metrics for data sub-circuits $\mathbf{c}^{D}$, we discard positional information and represent components as 2-tuples $(l,i)$.
@@ B.1 Prompt examples
Table˜2 presents prompt examples for each task in both input modalities.
@@
The amount of prompts per task is shown in Table˜3. The amount of prompts for the same task and modality varies between models due to tokenization considerations (e.g. all object colors in the spatial ordering task must be tokenized to a single token for positional alignment purposes—prompts that contain multi-token object colors are filtered out).
@@ B.2 Tasks Generation Details
For the textual variant, each prompt consists of two two-digit operands, ensuring positional alignment between different prompts (e.g., the second position always includes the singles digit of the first operand). This consistency is important because all analyzed models tokenize numbers into separate digits. For the visual variant, we create white-background images (7575×$338$ resolution) displaying the identical arithmetic calculations in large black font centered in the image, maintaining a simple presentation that focuses entirely on the calculation itself.
@@
For the textual variant, we select entities (to replace the [X] in the prompt, e.g. “Michael Jordan”) that, when combined with the prefix “Consider [X]”, result in exactly 5 tokens to ensure positional alignment. We use this number as it results in the highest number of possible entities in the given dataset. For the visual variant, we collect entity images with public licenses from Wikimedia Commons. We rank 10 potential images per entity using CLIP (Radford et al., 2021) scores and select the best match for each entity. We resize and center-crop each image (to 256256x256256 resolution). We manually review the images to remove any that contain “shortcut” hints (e.g., any image of a basketball player holding a basketball), to ensure models rely on factual recall rather than visual cues. In both variants, the counterfactual prompts for each template are sampled only from within the same template.
@@ B.3 Circuit Discovery and Evaluation
For our circuit discovery experiments (detailed in Section˜2 and Section˜4), we allocate each modality-specific task dataset with a 75/25 split: 75% of the prompts for discovery and 25% of the prompts for faithfulness evaluation. We divide prompts between the discovery and evaluation subsets such that all possible answers are distributed in the same ratio between the subsets.
When answers span multiple tokens (e.g. in the arithmetic task, where answers can be comprised of multiple digits), we measure the patching effect (Equation˜1) solely on the first token. We verify that the answers for each prompt and counterfactual prompt ($r$ and $r^{\prime}$) differ in the analyzed first token.
@@ B.4 Compute Resources
Our experiments were conducted on an NVIDIA L4040 node equipped with $8$ GPUs, each containing 4848GB of memory. Peak memory consumption occurred during circuit discovery operations on the Gemma-3-12B-Instruct model, that required the parallel use of $4$ GPUs. The complete experimental suite, including studies not featured in the final paper, consumed roughly 200–300 GPU hours.
@@ Appendix C Task Accuracies
In Table˜4 we report the accuracies of models on each of the tasks. We observe that across most tasks, VLMs achieve higher accuracy in the textual variant than the visual variant, motivating our analysis into this performance gap. This is compatible with earlier results on closed-source VLMs (Fu et al., 2024).
@@ Appendix D Additional Circuit Discovery Results
In Table˜5 we report the circuit size, in a percentage of the model’s component count. As described in Section˜4, the circuits are chosen to be the minimum-sized circuit with a faithfulness of over 80%. The faithfulness achieved by each circuit on each task and modality is reported in Table˜6.
@@
Furthermore, in Figure˜10, Figure˜11, Figure˜12, Figure˜13, Figure˜14 we present the patching importance scores for each model and task, summed per position and layer. These figures are complementary to Figure˜3.
@@ E.1 Iterative Back-patching
Extending the back-patching experiments in Section˜5, we explore if additional processing of visual data tokens yields further improvements. To do that, we perform back-patching iteratively (back-patching more than once). After identifying the optimal layers $l_{src},l_{dst}$ and layer window size for each model and task, we perform back-patching iteratively. Namely, after each patching of the window centered at $l_{dst}$, we continue the forward pass to calculate the window centered at $l_{src}$, and re-patch the window centered at $l_{dst}$. This process is repeated up to 10 times. As shown in Figure˜8, performance degrades after the first back-patching iteration across all but one task and model pairs, with each subsequent iteration further reducing model accuracy. We attribute this decline to representations becoming increasingly out-of-distribution with each iteration, making them progressively less compatible with the model’s learned parameters.
@@
E.2 Best Layers For Back-Patching
For each model and task, Table˜7 shows the values of $l_{src},l_{dst}$ and the layer window size that lead to the highest back-patching results, reported in Table˜1.
@@ E.2 Best Layers For Back-Patching
These optimal layers match our hypothesis across models: replacing visual data token representations at model layers when they are less textually-aligned, with more textually-aligned representations, yields the highest accuracy improvements. This pattern persists even in Gemma-3-12B, where visual data tokens demonstrate high alignment with textual analogs throughout most of the model (as shown in Figure˜7): tasks showing the greatest accuracy gains from back-patching (counting, arithmetic, and sentiment analysis, as detailed in Table˜1) benefit from patching in the earliest layers where the alignment of tokens with textual analogs is relatively low.
E.3 Back-Patching Control Experiments
@@
This section presents our back-patching control experiment results across models and tasks. The accuracy increases from back-patching visual data tokens (reported in Table˜1) can either stem from earlier textually-aligned representations or from additional compute. To isolate the effect of adding computational layers, we conduct the same back-patching analysis on analogous textual prompts. In this control experiment, the alignment with textual tokens is inherent, so any accuracy increase must stem from added compute. Across each back-patching setting—model, task and layer window size—we check for how many values of $l_{src}$ and $l_{dst}$ is the accuracy increase caused by back-patching larger than the control accuracy increase. Our findings (Figure˜9) show that in most settings, visual back-patching yields greater accuracy improvements than the control for a majority of layer values. A notable exception occurs in factual recall tasks for Qwen2-VL-7B and Gemma-3-12B, where back-patching shows minimal accuracy gains and is outperformed by the control in over half the cases. These results support our hypothesis that improved visual performance stems primarily from the early transition of visual tokens into textually-aligned representations, rather than from increased computational depth.
E.4 Back-Patching in General VQA
To demonstrate that back-patching generalizes beyond our constructed dataset in broader visual question-answering, we evaluate its impact on model accuracy using the VQAv2 (Goyal et al., 2017) and RealWorldQA (xAI, 2024) datasets, two common VQA benchmarks. We measure the accuracy of each VLM on each dataset without back-patching as a baseline, and measure it again when applying back-patching (as described in Section˜5). We present the results in Table˜8. For VQAv2, we measure the accuracy of each analyzed VLM on $3000$ randomly sampled visual prompts (ensuring no image repetition). For RealWorldQA, we measure the accuracy on the entire dataset of $765$ prompts. We limit the analysis to this amount of prompts due to computational requirements of exploring different back-patching settings. For both datasets, we also evaluate the mean and standard deviation of the accuracy by bootstrapping for $1000$ iterations with a full sample size. In both datasets, following our experimental setup for other tasks (Section˜B.3, we sample the model for a single forward pass, forcing an immediate answer. While this leads to generally lower results on the benchmarks, this setup is necessary due to the need to search for the best-performing back-patching settings, and ensures consistency across our experimental framework.
Table 8: Back-patching results on general VQA benchmarks. Back-patching leads to a significant increase in accuracy in general VQA benchmarks as well.
Dataset Setting Qwen2-7B-VL Pixtral-12B Gemma3-12B
VQAv2 Baseline $65.0\%$ $62.2\%$ $65.7\%$
Back-patching $8.2\%\pm 0.9\% 66$$6.2\%\pm 1.1\%$ $70.3\%\pm 1.3\%$
RealWorld Baseline $58.8\%$ $58.3\%$ $59.1\%$
Back-patching $61.6\%\pm 1.0\%$ $60.1\%\pm 1.2\%$ $60.0\%\pm 0.8\%$
The results confirm the average improvements observed in Section˜5. Since the visual prompts in VQAv2 and RealWorldQA do not have textual analogs, we do not perform the control experiment in this setting.
E.5 Back-patching across model scales
To evaluate if and how back-patching visual data representations affects accuracy across model scales, we perform back-patching on 3 scales of the PaliGemma2 VLM suite (Steiner et al., 2024): 3B, 10B and 28B parameters. We evaluate the models on our dataset of five tasks, and measure the mean accuracy and it’s standard deviation by bootstrapping for $1000$ iterations with full sample size. The baseline accuracies are presented in Table˜9. The back-patched accuracies are reported in Table˜10 and show that across most model scale and task pairs, back-patching increases model accuracy, by aligning visual token representations with their analogous textual representations earlier in the model.
Table 9: Baseline PaliGemma2 accuracies.
Model Size Counting Arithmetic Color Ordering Factual Recall Sentiment Analysis
3B $64.0\%$ $0.7\%$ $15.4\%$ $32.7\%$ $48.7\%$
10B $66.0\%$ $8.7\%$ $14.0\%$ $38.7\%$ $56.7\%$
28B $22.0\%$ $6.7\%$ $12.0\%$ $30.7\%$ $10.7\%$
Table 10: Back-patching accuracy increase across model scales. We evaluate how back-patching affects the accuracy of the PaliGemma2 VLM suite in three model sizes.
Model Size Counting Arithmetic Color Ordering Factual Recall Sentiment Analysis
3B $1.3\%\pm 2.9\% +$$31.6\%\pm 3.8\% +$$8.1\%\pm 3.4\% +$$6.5\%\pm 3.9\% +$$7.6\%\pm 3.9\%$
10B $3.4\%\pm 3.1\% +$$28.1\%\pm 4.1\% +$$11.9\%\pm 3.4\% +$$17.5\%\pm 4.0\% +$$16.1\%\pm 3.6\%$
28B $4.2\%\pm 3.6\% +$$24.1\%\pm 3.8\% +$$3.5\%\pm 2.9\% +$$5.3\%\pm 4.0\% +$$15.3\%\pm 3.6\%$
A note-worthy finding in the results is that the accuracy increase caused by back-patching is the largest in the 10B model size. We hypothesize this is caused by the overall lower performance of the 3B and the 28B model size. According to Steiner et al. (2024), the lower performance at the 28B size likely happens because this model scale is trained from scratch, as opposed to the smaller scales that are a result of model distillation.